# MFCLIP: Multi-modal Fine-grained CLIP for Generalizable Diffusion Face Forgery Detection

### v1

@@ Abstract
The rapid development of photo-realistic face generation methods has raised significant concerns in society and academia, highlighting the urgent need for robust and generalizable face forgery detection (FFD) techniques. Although existing approaches mainly capture face forgery patterns using image modality, other modalities like fine-grained noises and texts are not fully explored, which limits the generalization capability of the model. In addition, most FFD methods tend to identify facial images generated by GAN, but struggle to detect unseen diffusion-synthesized ones. To address the limitations, we aim to leverage the cutting-edge foundation model, contrastive language-image pre-training (CLIP), to achieve generalizable diffusion face forgery detection (DFFD). In this paper, we propose a novel multi-modal fine-grained CLIP (MFCLIP) model, which mines comprehensive and fine-grained forgery traces across image-noise modalities via language-guided face forgery representation learning, to facilitate the advancement of DFFD. Specifically, we devise a fine-grained language encoder (FLE) that extracts fine global language features from hierarchical text prompts. We design a multi-modal vision encoder (MVE) to capture global image forgery embeddings as well as fine-grained noise forgery patterns extracted from the richest patch, and integrate them to mine general visual forgery traces. Moreover, we build an innovative plug-and-play sample pair attention (SPA) method to emphasize relevant negative pairs and suppress irrelevant ones, allowing cross-modality sample pairs to conduct more flexible alignment. Extensive experiments and visualizations show that our model outperforms the state of the arts on different settings like cross-generator, cross-forgery, and cross-dataset evaluations.
@@ I Introduction
Photorealistic generation models like generative adversarial networks (GANs) [1, 2, 3] and denoising diffusion probabilistic models (DDPM) [4], have reached unprecedented progress in synthesizing highly realistic facial images. Thus, advancing face forgery detection (FFD) becomes a critical and urgent demand.
@@
Current FFD methods mainly involve three categories, spatial-based models [7, 8, 9, 10, 11], frequency-based models [12, 13, 14], and vision-language-based models [15, 16]. Spatial-based methods [7, 10] tend to mine domain-specific features in GAN-synthesized images, such as identity and background, which leads to poor generalization to unseen facial images generated by diffusion models. Frequency-based methods [17, 12, 18] aim to explore common face forgery traces in the frequency domain, since prior investigations [19, 20] have demonstrated that forgery artifacts produced by various synthesis approaches are mainly concentrated in the high-frequency domain, but they are inclined to capture high-frequency information at a coarse-grained level. As Fig. 1 (a) shows, face images synthesized by advanced diffusion-based models are so realistic that high-frequency noises are rarely observed. Only coarse-grained extraction of high-frequency noises is insufficient to dig comprehensive and universal forgery artifacts. Vision-language-based methods [15, 16] intend to study general face manipulation patterns using text-guided image forgery representation learning, but they struggle to introduce hierarchical fine-grained text prompts and achieve flexible cross-modal alignment, to facilitate the generalization to face images synthesized using diffusion models.
@@ II Related Work
Our research mainly involves two aspects: face forgery detection and vision-language models. In the following subsections, we discussed the two points, separately.
@@ II-A Face Forgery Detection
Existing efforts have achieved considerable progress in the field of FFD. Some methods focus on forgery artifacts in the spatial domain. Rossler et al.[22] employed an Xception pre-trained on ImageNet, to capture local spatial forgery artifacts. To explore comprehensive relations among image patches, a convolutional vision transformer (CViT) model [10] is designed to combine CNN with vision transformer (ViT) [8] for FFD, to mine global forgery traces in the spatial domain. Diffusion reconstruction error (DIRE) [9] utilizes the discrepancy between an input image and corresponding reconstruction for diffusion-generated image detection. In addition to employing spatial forgery patterns for FFD, there are also some frequency-aware approaches [12, 13, 14]. SFDG [14] is proposed to mine the relation-aware representations in spatial and frequency domains using dynamic graph learning. FreqNet [12] is devised to focus on high-frequency forgery traces across spatial and channel dimensions. To capture diverse forgery patterns across various modalities, TwoStream [13] is designed to integrate the high-frequency noise features with RGB content for generalized FFD. Recently, VLFFD [15] is proposed to acquire more generalization and interpretability for FFD, where fine-grained text prompts and coarse-grained original data are collaboratively employed to guide the coarse-and-fine co-training network learning. DD-VQA [16] introduces a multi-modal transformer framework, which boosts the learning of face forgery representations using a text and image contrastive learning scheme, to facilitate the FFD. By contrast, since there are significant discrepancies between authentic and forgery facial noise images extracted by SRM from the richest patches, we solely extract the SRM noises from the single richest patch, and combine them with global spatial forgery patterns to learn more general embeddings for DFFD. Furthermore, we boost common forged representations across image and noise modalities via fine-grained vision-language contrastive learning.
@@ II-B Vision-Language Models
Different from conventional image-based models that consist of an image feature encoder and a classifier to predict a fixed set of predefined sample categories, the vision-language models [23, 24, 25] such as contrastive language-image pre-training (CLIP) [21] simultaneously train an image encoder and a text encoder, to match image and text pairs from training datasets. CLIP facilitates inference on downstream tasks using a zero-shot linear classifier embedded with the class names or descriptions from the target dataset, showing robust image representations across various domains, including object detection and image generation. To circumvent the labour-intensive and inconsistent performance associated with prompt engineering for text encoders, CoOp [26] models the context words of a prompt using learnable vectors, and places the class token (i.e., names or descriptions) at the end or middle position. CFPL [27] is designed to study the various semantic prompts conditioned on different visual features for generalizable face anti-spoofing. Unlike traditional CLIP-based methods, our model is capable of capturing visual face forgery embeddings across image-noise modalities, and aligning the cross-modality features using the SPA method, adaptively.
@@ III-A Method Overview
To capture general multi-modal visual forgery patterns for generalizable DFFD, we design the multi-modal fine-grained CLIP model, to conduct the fine-grained text-guided visual forgery representation learning. As Fig. 3 shows, MFCLIP mainly consists of four key modules: fine-grained text generator (FTG), multi-modal vision encoder (MVE), fine-grained language encoder (FLE), and the sample pair attention (SPA) module. During the training phase, given a batch of input facial images $X$, MFCLIP generates the corresponding text prompts $T$ of each image from the GenFace dataset $D$ via the FTG module, and obtains the image-text pairs $(X,T)\in D$ with one-hot labels $y\in\big{\{}[0,1]^{T},[1,0]^{T}\}$. MFCLIP then captures the multi-modal visual forgery embeddings $X_{\text{v}}$ through the MVE module, and extracts abundant fine-grained language embeddings $T_{\text{l}}$ via the FLE module, respectively. Thereafter, $X_{\text{v}}$ and $T_{\text{l}}$ are transmitted to the SPA module to adaptively emphasize and suppress sample pairs, to generate $X_{\text{spa}}$. Meanwhile, $X_{\text{v}}$ is fed into a predictor to predict language features $T_{\text{l}}^{\text{pre}}$. Finally, the multilayer perceptron (MLP) head consisting of a full connection layer generates the final predictions $y_{\text{pre}}$. The kullback-leibler (KL) divergence loss function $\mathcal{L}_{\text{kl}}$ is leveraged to measure the difference between predicted language features $T_{\text{l}}^{\text{pre}}$ and the true language embeddings $T_{\text{l}}$. The cross-modality contrastive (CMC) loss $\mathcal{L}_{\text{cmc}}$ is used to align features between vision-language sample pairs. The cross-entropy (CE) loss $\mathcal{L}_{\text{ce}}$ is utilized to compute the discrepancy between the predicted label $y_{\text{pre}}$ and the ground truth label $y$. That is,
@@ III-C Multi-modal Vision Encoder
Unlike conventional multi-modal FFD methods [28, 17] that mainly concentrate on the interaction between the frequency or noise and RGB information at a coarse-grained level, we design the multi-modal vision encoder (MVE) to capture global spatial forgery traces as well as fine-grained noise manipulated patterns extracted from the richest patches, and integrate them in a simple and effective way, to achieve generalizable DFFD.
@@ III-E Sample Pair Attention
To enhance cross-modal feature alignment flexibly, we propose the sample pair attention (SPA) module. Unlike vanilla CLIP [21] which adjusts features using the static alignment mechanism, our MFCLIP dynamically emphasizes relevant sample pairs and suppresses irrelevant ones, enabling more flexible and effective cross-modal representation learning in a simple and efficient way.
@@ III-F Loss Function
Figure 6: Comparison of the performance of detectors on various generators. We averaged the AUC scores of the model listed in Table III on the images from the same generator.
TABLE I: Cross-forgery generalization. ACC and AUC scores (%) on remaining manipulations, after training using one manipulation.
@@
TABLE II: Cross-generator evaluation on FS. ACC and AUC scores (%) on remaining generators, after training using one generator.
@@ IV Experiments
To investigate the performance of our MFCLIP detector, we conducted experiments using five FFD datasets: faceforensics++ (FF++) [22], deepfake detection challenge (DFDC) [30], Celeb-DF [31], deeperforensics (DF-1.0) [32], and GenFace [33]. The remainder of this section is organized as follows: 1) the experiment setup including implementation details and datasets is introduced. 2) the performance assessments and comparisons based on five FFD datasets are outlined.
@@ IV-A Experiment Setup
Implementation details. We developed the detector using PyTorch on the Tesla V100 GPU with 32GB memory. The number of blocks $B$ and $L$ in MFCLIP is set to 3 and 6, respectively. The patch size $p$, feature dimension $d$, and batch size $b$ are set to 112, 512, and 24, respectively. We set the channel $c$, height $h$, and width $w$ to 64, 8, and 8, respectively. Our model is trained with the Adam optimizer [34] with a learning rate of 1e-4 and weight decay of 1e-3. We used the scheduler to drop the learning rate by ten times every 15 epochs. We used accuracy (ACC) and area under the receiver operating characteristic curve (AUC) as evaluation metrics.
Datasets. We employed the fine-grained face forgery dataset GenFace [33] for DFFD, where we conducted the cross-forgery protocol and cross-generator evaluation. We used four deepfake datasets to evaluate the generalization and robustness of networks: FF++ [22], DFDC [30], Celeb-DF [31], and DF-1.0 [32].
TABLE III: Cross-diffusion generalization. ACC and AUC scores (%) on remaining diffusion-based generators, after training using one diffusion-based generator.
@@ IV-B Comparison with the State of the Art
We evaluate the performance of state-of-the-art deepfake detectors on diffusion-generated images using GenFace. We selected various detectors such as CNN-based Xception [7] and FreqNet [12], transformer-based ViT [8], CViT [10] and FatFormer [29], vision-language-based CLIP [21], VLFFD [15] as well as DD-VQA [16], and DIRE [9] detecting diffusion-generated images.
@@
TABLE IV: Cross-generator evaluation on AM. ACC and AUC scores (%) on remaining generators, after training using one generator.
@@
TABLE V: Cross-dataset generalization. ACC and AUC scores on FF++, Celeb-DF, DFDC, and DF-1.0 after training using FF++.
@@
Cross-forgery evaluation. To investigate the generalization of various detectors, we performed cross-forgery tests. We trained models using images of one manipulation, and tested them on those of remaining manipulations. As Table I shows, the performance of our methods outperforms most detectors, demonstrating the superior generalization of MFCLIP. Specifically, for the vision-language-based models, the AUC of our model is nearly 26.80%, 18.78%, and 14.66% higher than that of CLIP, VLFFD, and DD-VQA, respectively, on EFS after training using AM, which is attributed to the powerful learning capabilities of our model for fine-grained text-guided multi-modal face forgery representations.

### v2

@@ Abstract
The rapid development of photo-realistic face generation methods has raised significant concerns in society and academia, highlighting the urgent need for robust and generalizable face forgery detection (FFD) techniques. Although existing approaches mainly capture face forgery patterns using image modality, other modalities like fine-grained noises and texts are not fully explored, which limits the generalization capability of the model. In addition, most FFD methods tend to identify facial images generated by GAN, but struggle to detect unseen diffusion-synthesized ones. To address the limitations, we aim to leverage the cutting-edge foundation model, contrastive language-image pre-training (CLIP), to achieve generalizable diffusion face forgery detection (DFFD). In this paper, we propose a novel multi-modal fine-grained CLIP (MFCLIP) model, which mines comprehensive and fine-grained forgery traces across image-noise modalities via language-guided face forgery representation learning, to facilitate the advancement of DFFD. Specifically, we devise a fine-grained language encoder (FLE) that extracts fine global language features from hierarchical text prompts. We design a multi-modal vision encoder (MVE) to capture global image forgery embeddings as well as fine-grained noise forgery patterns extracted from the richest patch, and integrate them to mine general visual forgery traces. Moreover, we build an innovative plug-and-play sample pair attention (SPA) method to emphasize relevant negative pairs and suppress irrelevant ones, allowing cross-modality sample pairs to conduct more flexible alignment. Extensive experiments and visualizations show that our model outperforms the state of the arts on different settings like cross-generator, cross-forgery, and cross-dataset evaluations. Our code will be available at https://github.com/Jenine-321/MFCLIP.
@@ I Introduction
Photorealistic generation models like variational autoencoder (VAE) [1], generative adversarial networks (GANs) [2, 3, 4], denoising diffusion probabilistic models (DDPM) [5], autoregressive networks [6, 7], and flow-based methods [8] have reached unprecedented progress in synthesizing highly realistic facial images. Thus, advancing face forgery detection (FFD) becomes a critical and urgent demand.
@@
Current FFD methods mainly involve three categories, spatial-based models [11, 12, 13, 14, 15, 16], frequency-based models [17, 18, 19], and vision-language-based models [20, 21]. Spatial-based methods [11, 14] tend to mine domain-specific features in GAN-synthesized images, such as identity and background, which leads to poor generalization to unseen facial images generated by diffusion models. Frequency-based methods [22, 17, 23] aim to explore common face forgery traces in the frequency domain, since prior investigations [24, 25] have demonstrated that forgery artifacts produced by various synthesis approaches are mainly concentrated in the high-frequency domain, but they are inclined to capture high-frequency information at a coarse-grained level. As Fig. 1 (a) shows, face images synthesized by advanced diffusion-based models are so realistic that high-frequency noises are rarely observed. Only coarse-grained extraction of high-frequency noises is insufficient to dig comprehensive and universal forgery artifacts. Vision-language-based methods [20, 21] intend to study general face manipulation patterns using text-guided image forgery representation learning, but they struggle to introduce hierarchical fine-grained text prompts and achieve flexible cross-modal alignment, to facilitate the generalization to face images synthesized using diffusion models.
@@ II Related Work
Our research mainly involves two aspects: face forgery detection and vision-language models. In the following subsections, we discuss the two points, separately.
@@ II-A Face Forgery Detection
Existing efforts have achieved considerable progress in the field of FFD [27]. Some methods [28, 29] focus on forgery artifacts in the spatial domain. Rossler et al. [30] employ an Xception pre-trained on ImageNet, to capture local spatial forgery artifacts. To explore comprehensive relations among image patches, a convolutional vision transformer (CViT) model [14] is designed to combine CNN with vision transformer (ViT) [12] for FFD, to mine global forgery traces in the spatial domain. Diffusion reconstruction error (DIRE) [13] utilizes the discrepancy between an input image and corresponding reconstruction for diffusion-generated image detection. In addition to employing spatial forgery patterns for FFD, there are also some frequency-aware approaches [17, 19]. Hsu et al. [31] propose a diffusion-based image inpainting method that combines stochastic sampling with feature keypoint guidance, to improve small-area fingerprint recognition. SFDG [19] is proposed to mine the relation-aware representations in spatial and frequency domains using dynamic graph learning. FreqNet [17] is devised to focus on high-frequency forgery traces across spatial and channel dimensions. To capture diverse forgery patterns across various modalities, TwoStream [18] is designed to integrate the high-frequency noise features with RGB content for generalized FFD. Recently, VLFFD [20] is proposed to achieve generalization and interpretability for FFD, where fine-grained text prompts and coarse-grained original data are collaboratively employed to guide the coarse-and-fine co-training network learning. DD-VQA [21] introduces a multi-modal transformer, which boosts the learning of face forgery representations using a text and image contrastive learning scheme. By contrast, since there are significant discrepancies between authentic and forgery facial noise images extracted by SRM from the richest patches, we solely extract the SRM noises from the richest patch, and combine them with global spatial forgery patterns to learn more general embeddings. Furthermore, we boost common forged representations across image and noise modalities via fine-grained vision-language matching.
@@ II-B Vision-Language Models
Different from conventional image-based models that consist of an image feature encoder and a classifier to predict a fixed set of predefined sample categories, the vision-language models [32, 26] such as contrastive language-image pre-training (CLIP) [26] simultaneously train an image encoder and a text encoder, to match image and text pairs from training datasets. CLIP facilitates inference on downstream tasks using a zero-shot linear classifier embedded with the class names or descriptions from the target dataset. To circumvent the labour-intensive and inconsistent performance associated with prompt engineering for text encoders, CoOp [33] models the context words of a prompt using learnable vectors, and places the class token (i.e., names or descriptions) at the end or middle position. CFPL [34] is designed to study the various semantic prompts conditioned on different visual features for generalizable face anti-spoofing. Unlike traditional CLIP-based methods, our model is capable of capturing visual face forgery embeddings across image-noise modalities, and aligning the cross-modality features using the SPA method, adaptively.
@@ III-A Method Overview
To capture general multi-modal visual forgery patterns for generalizable DFFD, we design the multi-modal fine-grained CLIP model, to conduct the fine-grained text-guided visual forgery representation learning. Different from vision-language-based methods that focus on image and text modalities, our model explores fine-grained noise, image, and text modalities. As Fig. 3 shows, MFCLIP mainly consists of four key modules: fine-grained text generator (FTG), multi-modal vision encoder (MVE), fine-grained language encoder (FLE), and the sample pair attention (SPA) module. During the training phase, given a batch of input facial images $X$, MFCLIP generates the corresponding text prompts $T$ of each image from the GenFace dataset $D$ via the FTG module, and obtains the image-text pairs $(X,T)\in D$ with one-hot labels $y\in\big{\{}[0,1]^{T},[1,0]^{T}\}$. MFCLIP then captures the multi-modal visual forgery embeddings $X_{\text{v}}$ through the MVE module, and extracts abundant fine-grained language embeddings $T_{\text{l}}$ via the FLE module, respectively. Thereafter, $X_{\text{v}}$ and $T_{\text{l}}$ are transmitted to the SPA module to adaptively emphasize and suppress sample pairs, to generate $X_{\text{spa}}$. Meanwhile, $X_{\text{v}}$ is fed into a predictor with a fully connected layer to predict language features $T_{\text{l}}^{\text{pre}}$. Finally, the multilayer perceptron (MLP) head consisting of a full connection layer generates the final predictions $y_{\text{pre}}$. The kullback-leibler (KL) divergence loss function $\mathcal{L}_{\text{kl}}$ is leveraged to measure the difference between predicted language features $T_{\text{l}}^{\text{pre}}$ and the true language embeddings $T_{\text{l}}$. The cross-modality contrastive (CMC) loss $\mathcal{L}_{\text{cmc}}$ is used to align features between vision-language sample pairs. The cross-entropy (CE) loss $\mathcal{L}_{\text{ce}}$ is utilized to compute the discrepancy between the predicted label $y_{\text{pre}}$ and the ground truth label $y$. That is,
@@ III-C Multi-modal Vision Encoder
Unlike multi-modal FFD models [35, 22] that mainly concentrate on the interaction between the frequency or noise and RGB information at a coarse-grained level, we design the multi-modal vision encoder (MVE) to capture global image forgery traces as well as fine-grained noise patterns extracted from the richest patches, and integrate them simply and effectively.
@@ III-E Sample Pair Attention
To enhance cross-modal feature alignment flexibly, we propose the sample pair attention (SPA) module. Unlike traditional vision-language-based models like vanilla CLIP [26], VLFFD [20], and DD-VQA [21], which adjust features using the static alignment mechanism, MFCLIP dynamically emphasizes relevant sample pairs and suppresses irrelevant ones, enabling more flexible and effective cross-modal representation learning simply and efficiently.
@@ III-F Loss Function
The total loss is defined as follows:
$\mathcal{L}=\mathcal{L}_{\text{cmc}}+\mathcal{L}_{\text{ce}}+\mathcal{L}_{\text{kl}}$ (15)
TABLE I: Cross-forgery generalization. ACC and AUC scores (%) on remaining manipulations, after training using one manipulation in GenFace. EFS represents entire face synthesis, AM means attribute manipulation, and FS denotes face swapping.
@@
TABLE II: Cross-generator evaluation on FS. ACC and AUC scores (%) on remaining generators, after training using one generator in GenFace. FS denotes face swapping.
@@ IV Experiments
To investigate the performance of our MFCLIP detector, we conduct experiments using five FFD datasets: faceforensics++ (FF++) [30], deepfake detection challenge (DFDC) [37], Celeb-DF [38], deeperforensics (DF-1.0) [39], and GenFace [40]. The remainder of this section is organized as follows: 1) the experiment setup including implementation details and datasets is introduced. 2) the performance assessments and comparisons based on five FFD datasets are outlined.
@@ IV-A Experiment Setup
IV-A1 Implementation details
We develop the detector using PyTorch on a Tesla V100 GPU with 32GB memory. The time it takes to train MFCLIP is approximately 30 hours, and the model is trained for 30 epochs. The number of blocks $B$ and $L$ in MFCLIP is set to 3 and 6, respectively. The patch size $p$, feature dimension $d$, and batch size $b$ are set to 112, 512, and 24, respectively. We set the channel $c$, height $h$, and width $w$ to 64, 8, and 8, respectively. Our model is trained with the Adam optimizer [41] with a learning rate of 1e-4 and weight decay of 1e-3. We use the scheduler to drop the learning rate by ten times every 15 epochs. We use accuracy (ACC) and area under the receiver operating characteristic curve (AUC) as evaluation metrics. We measure fréchet inception distance (FID) [42] to assess the quality of generated face images, by comparing the feature distributions of real and fake images. The lower FID scores indicate higher similarity and better image quality.
IV-A2 Datasets
We use the fine-grained face forgery dataset GenFace [40] for DFFD, where we conduct the cross-generator and cross-forgery evaluation. We use four deepfake datasets to evaluate the generalization and robustness of networks: FF++ [30], DFDC [37], Celeb-DF [38], and DF-1.0 [39].
GenFace. GenFace contains a large number of facial images generated by different generators, which mainly involve three forgeries: entire face synthesis (EFS), attribute manipulation (AM), and face swapping (FS). EFS proposes to create non-existing forgery faces from random noises using generative methods. AM intends to manipulate the facial attributes of a real image to yield a forgery face. FS refers to swapping the identity of source images with that of target images. Each forgery includes several specific generators. Specifically, EFS comprises diffusion-based generators such as DDPM [5], LatDiff [43], CollDiff [44], and GAN-based ones like StyleGAN2 [4] and StyleGAN3 [3]. AM embraces diffusion-based generators including Diffae [45], as well as GAN-based ones such as LatTrans [46], IAFaces [47], etc. FS incorporates diffusion-based generators such as DiffFace [48], and GAN-based ones like FSLSD [49] and FaceSwapper [50]. The face image is scaled to a size of 224$\times$224. We conduct the normalization by dividing the image pixel values by 255, without data augmentation operations.
FF++. FF++ is a widely used benchmark dataset for FFD, which consists of videos with varying compression levels, including raw, high quality (HQ), and low quality (LQ). We use the HQ videos along with the official splits: 720 videos for training, 140 videos for validation, and 140 videos for testing. We take 20 frames at fixed intervals for each video. We leverage dlib [51] to extract square facial regions, and then resize them to a fixed resolution of 224$\times$224 pixels. We conduct the normalization by dividing the image pixel values by 255, without data augmentation operations.
DFDC. DFDC is a large-scale dataset containing over 100,000 videos with diverse manipulation techniques and realistic distortions. The data preprocessing way of DFDC is the same as that of FF++.
Celeb-DF. Celeb-DF is a high-quality deepfake dataset which contains 590 real videos and 5,639 deepfake ones. The data preprocessing way of Celeb-DF is the same as that of FF++.
DF-1.0. DF-1.0 is a large-scale face forgery dataset, comprising 60,000 videos with a total of 17.6 million frames. Various real-world perturbations are used to attain a more challenging benchmark with larger scale and higher diversity. The data preprocessing way of DF-1.0 is the same as that of FF++.
TABLE III: Cross-diffusion generalization. ACC and AUC scores (%) on remaining diffusion-based generators, after training using one diffusion-based generator in GenFace.
@@ IV-B Comparison with the State of the Art
We evaluate the performance of state-of-the-art deepfake detectors on diffusion-generated images using GenFace. We select various detectors such as CNN-based (i.e., Xception [11] and FreqNet [17]) , transformer-based (i.e., ViT [12], CViT [14], and FatFormer [36]) , vision-language-based models (i.e., CLIP [26], VLFFD [20], and DD-VQA [21]), and methods dedicated to detecting diffusion forgery images. (e.g., DIRE [13] and CAEL [40]). To ensure a fair comparison, we implement the SOTA methods using publicly available codebases and follow the original papers’ configurations. All models are trained and evaluated using the same dataset settings. The specific implementation details are as follows:
Xception and FreqNet. Since the Xception and FreqNet model codes have been open-sourced, to ensure a fair comparison, we train and test the open-source models using hyperparameter settings in the original paper, and the same dataset protocol as MFCLIP, until convergence, respectively.
CViT and FatFormer. We train and test the open-source CViT and FatFormer model using hyperparameter settings in the original paper, and the same dataset protocol as MFCLIP, until convergence, respectively.
ViT and CLIP. CLIP tends to be regarded as a foundational vision-language model that serves as a baseline. We use the official OpenAI implementation with the ViT-B/32 backbone. Pre-trained weights are downloaded from the official repository. We train CLIP on our datasets using the Adam optimizer with a learning rate of 1e-5 and weight decay of 1e-4, until convergence. The open-source ViT-B model is trained from scratch on our datasets via the Adam optimizer with a learning rate of 1e-4 and weight decay of 1e-3, until convergence.
VLFFD and DD-VQA. Since VLFFD and DD-VQA do not provide open-source implementations, we follow the descriptions in the original papers, to replicate the models using PyTorch frameworks. To ensure the correctness of our implementation, we compare the performance of replicated models with the results reported in the original papers on standard benchmarks. To ensure a fair comparison, we train and test the replicated model using hyperparameter settings in the original paper, and the same dataset protocol as MFCLIP, until convergence.
DIRE and CAEL. DIRE leverages the discrepancy between an input image and the corresponding reconstruction for diffusion image detection. CAEL integrates the discriminative edges global features with appearance global manipulated representations, to detect diffusion face forgery images. We train and test the CAEL and open-source DIRE model using hyperparameter settings in the original paper, and the same dataset protocol as MFCLIP, until convergence, respectively.
@@
TABLE IV: Cross-generator evaluation on AM. ACC and AUC scores (%) on remaining generators, after training using one generator in GenFace.
@@
TABLE V: Cross-dataset generalization. * denotes the results of the models are reproduced by ourselves. For the remaining models, we use the results reported in the paper. ACC and AUC scores on FF++, Celeb-DF, DFDC, and DF-1.0 after training using FF++.
@@
IV-B1 Cross-generator evaluation